# Coupled Lindblad pseudomode theory for simulating open quantum systems

### v1

@@ Abstract
Lorentzian pseudomode [4] $\mathrm{poly}(T/\varepsilon)$
@@
Introduction.– Open quantum systems, which describe quantum systems interacting with their environment, play a fundamental role in various fields, including quantum optics, condensed matter physics, chemical physics, and quantum information science [13, 14, 15, 16]. A common setting involves a system that is linearly coupled to a Gaussian environment with continuous degrees of freedom (e.g., over the frequency $\omega$). Upon tracing out the environment, the resulting system dynamics are generally non-Markovian. When simulating the non-Markovian dynamics on a classical or quantum computer via an approximate representation, two key questions arise: (1) Efficiency: Can the simulation be performed using minimal computational resources? (2) Physicality: Does the approximate dynamics correspond to a physically realizable process?
We focus on approximation schemes that employ a finite number of auxiliary bath modes (often called bath, environment, discrete mode, or pseudomode in various settings). The efficiency of such schemes is characterized by the number of modes required to capture the environment. The goal is to simulate all bounded observables on the system up to time $T$ within precision $\varepsilon$. This task is often reduced to the problem of fitting the bath correlation function (BCF) up to time $T$ with comparable precision [17, 18, 19, 20, 21]. An efficient approximation scheme means that the number of modes scales only as $\text{polylog}(T/\varepsilon)$.
By physicality, we mean that the joint dynamics of the system and auxiliary bath can be realized as a valid physical process. In particular, the dynamics can be implemented as a quantum channel, ensuring that the resulting solution map is completely positive and trace-preserving (CPTP). The CPTP property, in turn, guarantees numerical stability when simulating the dynamics on a classical computer. Physicality is also essential for efficient implementation on a quantum computer [22, 23, 24, 25, 26].
Table 1 summarizes and compares existing schemes. One of the earliest and most widely used approaches is the unitary discrete mode representation, where both the system and auxiliary modes evolve under unitary dynamics [1, 2]. Despite its wide usage, the number of modes scales linearly in the simulation time $T$ [2, 27, 3]. Intuitively, this limitation arises because unitary dynamics of a finite system always lack dissipation, making it incapable of accurately modeling BCFs that decay over time.
@@
The pseudomode theory [28, 29, 17, 30, 9, 31, 32, 33, 34] aims at addressing this limitation by introducing dissipation in order to model energy relaxation correctly. In Refs. [28, 29, 17], the environment is represented by bath modes, each subject to Lindblad dissipation. The spectral density becomes a sum of Lorentzians, which can be referred to as a Lorentzian pseudomode, and its dynamics can be realized using a quantum channel. However, a fundamental drawback is that the tail of a Lorentzian exhibits only inverse polynomial rather than exponential decay in the frequency domain. Theoretical analyses indicate that even for smooth spectral densities, the number of pseudomodes scales as $\text{poly}(T/\varepsilon)$ [4], though the preconstant can be smaller than that of unitary discrete modes.
@@
Recent advances in pseudomode methods, including non-Hermitian [5, 6, 35, 36, 37] and quasi-Lindblad pseudomodes [7, 8], have significantly improved the efficiency. Under certain analytic conditions on the spectral density, the number of required pseudomodes scales as $\mathrm{polylog}(T/\varepsilon)$ [19, 8]. However, these dynamics are not completely positive (CP) and cannot be realized as quantum channels.
@@
All the finite mode approximations discussed so far are decoupled, meaning that the modes interact with the system but not with one another. There is another class of approximation schemes that allows couplings between pseudomodes while preserving the Lindblad form, and hence the resulting dynamics can still be realized as a quantum channel [9, 10, 11, 12]. Although introduced in different contexts, we collectively refer to these as coupled Lindblad pseudomode theory. In particular, empirical studies in Refs. [9, 10] demonstrated that only a small number of such coupled pseudomodes are sufficient to accurately approximate the BCF. However, it remains unclear whether this improvement is asymptotic or due to a smaller preconstant. Moreover, the construction in [9, 10, 11, 12] relies on non-convex optimization, which can be challenging to perform in practice.
@@
The contributions of this Letter are twofold: (1) We establish a direct connection between the coupled Lindblad pseudomode theory and the quasi-Lindblad pseudomode theory. Combined with the theoretical results of Ref. [8] on spectral density fitting, this provides a theoretical justification that the optimal parameterization cost of coupled Lindblad pseudomodes scales as $\mathrm{polylog}(T/\varepsilon)$. (2) Inspired by the realization problem in control theory [38], we develop a robust numerical algorithm for constructing the coupled modes that avoids the non-convex optimization used in Refs. [9, 10, 11, 12]. This can significantly simplify the process of finding coupled Lindblad pseudomodes and lead to more accurate pseudomodes without increasing their number. Taken together, this provides a firm theoretical and computational foundation for the widespread application of this theory.
@@
Model setup.– For concreteness, we consider a spin-boson Hamiltonian $\hat{H}=\hat{H}_{\text{S}}+\hat{H}_{\text{B}}+\hat{H}_{\text{SB}}$, where $\hat{H}_{\text{B}}=\int_{0}^{\infty}\omega\hat{b}_{\omega}^{\dagger}\hat{b}_{\omega}\mathrm{d}\omega$, $\hat{H}_{\text{SB}}=\hat{S}\hat{B}$, $\hat{S},\hat{B}$ are Hermitian, and $\hat{B}=\int_{0}^{\infty}\sqrt{J(\omega)}(\hat{b}_{\omega}+\hat{b}_{\omega}^{\dagger})\mathrm{d}\omega$. Here, $\hat{b}_{\omega}$ is a bosonic annihilation operator and $J(\omega)$ is the spectral density. Assuming an initially factorized state $\hat{\rho}(0)=\hat{\rho}_{\text{S}}(0)\otimes\hat{\rho}_{\text{B}}(0)$ where $\hat{\rho}_{\text{B}}(0)$ is assumed to be the equilibrium state corresponds to $\hat{H}_{\text{B}}$, our goal is to compute the dynamics of the system-reduced density operator $\hat{\rho}_{\text{S}}(t)=\operatorname{tr}_{\text{B}}\left({\mathrm{e}}^{-\mathrm{i}\hat{H}t}\hat{\rho}(0){\mathrm{e}}^{\mathrm{i}\hat{H}t}\right)$, where the influence of environment is fully captured by the bath correlation function (BCF),
@@
Pseudomode theory.– Pseudomode theory introduces a set of finite number of auxiliary modes (denoted as A). If the pseudomode BCF matches the original BCF for all $t\in[0,T]$, then the corresponding pseudomode dynamics exactly reproduces the reduced system density matrix $\hat{\rho}_{\text{S}}(t)$ up to time $T$ [17].
The dynamics of the coupled Lindblad pseudomode are given as follows:
@@
with the bath initially in the vacuum state, i.e., $\hat{\rho}_{\text{A}}(0)=\mathinner{|{\boldsymbol{0}}\rangle}\!\!\mathinner{\langle{\boldsymbol{0}}|}$. Here, $H=H^{\dagger}$, $\Gamma\succeq 0$, and $\hat{A}=\sum_{k}g_{k}\hat{b}_{k}+\overline{g_{k}}\hat{b}_{k}^{\dagger}$, where we denote a positive semidefinite (definite) matrix $M$ as $M\succeq 0$ ($M\succ 0$). The conditions on $H$ and $\Gamma$ ensure that the dynamics of $\hat{\rho}_{\text{SA}}^{\text{c}}$ are CPTP, i.e. physical. The term coupled modes indicates that both $H$ and $\Gamma$ can be dense matrices, meaning they could contain off-diagonal elements that mediate couplings between modes. In contrast, the Lorentzian pseudomode theory assumes $H$ and $\Gamma$ are diagonal, meaning that the modes are decoupled. The BCF of the coupled Lindblad pseudomode, denoted as $C^{\text{c}}(t)=\operatorname{tr}\left(\hat{A}(t)\hat{A}(0)\hat{\rho}_{\text{A}}(0)\right)$, is given as follows [21, 10]:
@@
Another pseudomode theory we will compare with is the quasi-Lindblad theory [7]. It includes additional system-bath dissipation to Eq. (2),
@@
The system-bath dissipation $\boldsymbol{D}_{\text{SA}}$ in general breaks the CP condition when there is no dissipation acting on the system. Violating the CP condition can induce instabilities in the quasi-Lindblad dynamics, posing challenges for classical simulation  [^1] There is a subtle concept called the Hamiltonian-induced stability [7], which may stabilize the quasi-Lindblad dynamics, even under the CP violation, but in the worst case, the quasi-Lindblad dynamics show the instability.. We note that the hierarchical equations of motion (HEOM) approach [43, 39] may also encounter similar stability challenges [44, 45, 46].
In the context of quantum simulation, the loss of the CP condition indicates that the dynamics can no longer be efficiently  [^2] To encode the evolution using a quantum channel, an exponentially large subnormalization factor in $T$ must be introduced, rendering long-time simulation on a quantum computer prohibitively expensive. implemented as a quantum channel. The coupled Lindblad dynamics, however, are inherently CP and TP, ensuring numerical stability and compatibility with quantum hardware. The coupling between bosonic modes can also be realized in analog ion-trap-based quantum simulators [25, 48, 49, 26].
Efficient construction of coupled modes with $\mathrm{polylog}(T/\varepsilon)$ scaling.– It may seem that the coupled Lindblad dynamics is highly overparameterized, as it involves $\mathcal{O}(N^{2})$ parameters compared to only $\mathcal{O}(N)$ in the decoupled finite-mode approximation. However, the positivity condition $\Gamma\succeq 0$ is nontrivial to satisfy and imposes strong constraints on the parameter space. It is not clear a priori whether the coupled Lindblad pseudomode formulation can significantly outperform the Lorentzian Lindblad model, even with the additional degrees of freedom. Moreover, the increased number of parameters makes the model more difficult to optimize using black-box non-convex optimization solvers.
We first discuss the connection between the coupled Lindblad BCF $C^{\text{c}}(t)=g^{\dagger}\mathrm{e}^{-\mathrm{i}Kt}g$ Eq. (3) and the quasi-Lindblad BCF $C^{\text{q}}(t)=l^{\dagger}\mathrm{e}^{-\mathrm{i}\Lambda t}r$ Eq. (5). The result is summarized in the following theorem. We will then describe an algorithm for finding the parameters in the coupled Lindblad pseudomode formulation that avoids non-convex optimization.
@@
[…]
@@
The first part of the theorem parallels the result of Ref. [17], which establishes that, for fixed $\hat{H}_{\text{S}}$, $\hat{S}$, and $\hat{\rho}_{\text{S}}(0)$, the reduced system dynamics are uniquely determined by the BCF. Therefore, we focus on the second part. A gauge transformation, $\Lambda\rightarrow K=X\Lambda X^{-1}$, $l^{\dagger}\rightarrow l^{\dagger}X^{-1}$, $r\rightarrow Xr$, with an invertible matrix $X$, leaves $C^{\text{q}}(t)$ invariant. The gauge-transformed BCF takes the coupled Lindblad form if the following conditions hold: (a) $g=(l^{\dagger}X^{-1})^{\dagger}=Xr$, (b) $\Gamma=(K^{\dagger}-K)/2\mathrm{i}\succeq 0$. These two conditions correspond to the Hermiticity and the positivity of the dynamics, respectively. Introducing $Y=X^{\dagger}X$, we rewrite these as the equality and inequality constraints in Eq. 7 by multiplying $X$ and $X^{{\dagger}}$ to conditions (a) and (b). In addition, $X$ being invertible indicates that $Y\succ 0$, which makes conditions (a) and (b) equivalent to Eq. 7. ∎
Eq. 7 implies that when the feasibility condition is satisfied, the number of coupled Lindblad pseudomodes is no greater than that of quasi-Lindblad pseudomodes. Consequently, the $\mathrm{polylog}(T/\varepsilon)$ scaling of the latter extends directly to the coupled Lindblad setting (see Table 1).
In practice, the feasibility condition Eq. (7) may not hold exactly. Nonetheless, we construct a numerical procedure that minimally violates it by solving the following least-squares problem with semidefinite constraints:
@@
This problem could be solved efficiently via a semidefinite programming (SDP) solver. Setting $X=\sqrt{Y}$, we recover the parameters $g$, $H$, and $\Gamma$ in Eq. (3). This approach avoids the non-convex optimization used in prior works [10, 9, 11, 12].
It is worth noting that the fitting of $C^{\text{q}}(t)$ in Eq. 5 is often performed in the time domain, using signal processing algorithms such as ESPRIT [50, 7] and Prony algorithm [51]. However, if the bath information is provided in the frequency domain rather than in the time domain, a Fourier transform of the spectral density is required, which can introduce additional approximation errors, particularly when $J(\omega)$ is only available on a discrete frequency grid with limited accuracy.
In what follows, we propose a new procedure for obtaining the parameters in the coupled Lindblad theory directly from the BCF in the frequency domain, given by $\widetilde{C}(\omega)=\text{sign}(\omega)J(|\omega|)/(1-e^{-\beta\omega})$ in the spin-boson model with an inverse temperature $\beta$. Our algorithm is motivated by the realization problem in control theory [52, 38], which aims to fit a given scalar function $f$ in the form $f(\omega)=l^{\dagger}(K-\omega I)^{-1}r$. The algorithm is highly robust and only requires applying a singular value decomposition (SVD) to a certain Loewner matrix [38] constructed from the sampled data $f(\omega)$.
$\widetilde{C}(\omega)\approx\mathrm{Im}(g^{\dagger}(K-\omega I)^{-1}g),\quad K=H-\mathrm{i}\Gamma.$ (9)
@@
Figure 1: (a) For a fixed precision $\varepsilon=10^{-6}$ in fitting $C(t)$ for $t\in[0,T]$, we plot the number of modes required, $N$, against the maximum simulation time $T$. The numbers of coupled Lindblad and quasi-Lindblad pseudomodes scale as $\mathcal{O}(\log T)$ in contrast to the $\mathcal{O}(T)$ scaling in the unitary and Lorentzian modes. (b) For a fixed $T=10$, we plot $\varepsilon$ versus $N$, where the coupled Lindblad and quasi-Lindblad methods achieve a significantly faster convergence rate.
@@
Numerical results.– First, we verify the effectiveness of fitting the BCF $C(t)$ using the coupled Lindblad pseudomode theory by comparing it with other pseudomode approaches. Our target is to fit $C(t)$ for $t\in[0,T]$, derived from the Ohmic spectral density $J(\omega)=\omega\mathrm{e}^{-\omega/\omega_{c}}$ for $\omega\geq 0$ at zero temperature and $\omega_{c}=1$. For the coupled Lindblad fitting, we first fit $\widetilde{C}(\omega)$ in the frequency domain using the realization-based method, and further refine the result by using it as an initial guess for a gradient-based optimization of $C(t)$ in the time domain. We illustrate $T$-dependence (Fig. 1(a)) and $\varepsilon$-dependence (Fig. 1(b)) of the number of modes $N$, where we use the averaged $L^{2}$ error $\varepsilon$ [53, Eq. S8] as a measure of precision. The results of Fig. 1(a) with a fixed precision $\varepsilon=10^{-6}$ confirm that the coupled Lindblad pseudomode exhibits $N=\mathcal{O}(\log T)$ scaling, similar to the quasi-Lindblad pseudomode [8], in contrast to $N=\mathcal{O}(T)$ scaling in the unitary and Lorentzian modes. In Fig. 1(b), with a fixed $T=10$, all methods exhibit $N=\mathcal{O}(\log(1/\varepsilon))$ scaling, but the coupled Lindblad and quasi-Lindblad approaches converge significantly faster. The performance of the coupled Lindblad pseudomode theory closely matches that of the quasi-Lindblad pseudomode, indicating that the violation of the feasibility condition is small.
@@
Next, we demonstrate that the coupled Lindblad pseudomode theory can accurately capture real-time system dynamics. In Fig. 2(a), we present the population dynamics of the spin-boson model with $\hat{H}_{\text{S}}=\frac{\omega_{e}}{2}\hat{\sigma}_{z}$ and $\hat{S}=\hat{\sigma}_{x}$, where $\hat{\sigma}_{z}$ and $\hat{\sigma}_{x}$ are the Pauli operators. We follow the setup in Ref. [10] by choosing the Lorentzian-like spectral density, $J(\omega)=\frac{2g^{2}\kappa\omega_{c}\omega/\pi}{(\omega_{c}^{2}-\omega^{2})^{2}+\kappa^{2}\omega^{2}}$ for $\omega\geq 0$ and $J(\omega)=0$ for $\omega<0$. We focus on the ultra-strong coupling regime, with parameters $\omega_{c}=\omega_{e}=0.58,\ g=0.25,$ and $\kappa=0.1\text{meV}$ [10]. Fig. 2(a) describes population dynamics $n_{0}(t)=\langle 0|\hat{\rho}_{S}(t)|0\rangle$ evolved from the initial state $\hat{\rho}_{S}(0)=|0\rangle\!\langle 0|$. We use $N=2$ and $N=4$ coupled pseudomodes in this simulation, compared to the reference data obtained from unitary dynamics with a large discretization $N=400$. A comparison with the pseudomode dynamics in Ref. [10] for $N=10$ shows that our approach achieves similar accuracy with only $N=4$ modes. The pseudomode fitting in Ref. [10] is based on a non-convex optimization of $J(\omega)$, with a penalty applied to $J(\omega)$ for $\omega<0$ to suppress unphysical contributions associated with the artificial pumping at negative frequencies. Nevertheless, as shown in Fig. 2(b), even without such explicit penalization, the fitted BCF already exhibits a very small contribution (on the order of $10^{-5}$) in the negative frequency region with as few as $N=4$ modes, and this contribution becomes negligible (around $10^{-9}$) when using $N=10$ modes. This highlights the robustness of our convex optimization strategy, which constructs the coupled modes even with sharp changes in $J(\omega)$.
@@
Figure 3: Normalized absorption spectrum $S(\omega)$ for the dimer model with two different environments $J_{0}(\omega)$ (left column) and $J_{1}(\omega)$ (right column), at zero temperature (first row) and finite temperature (77K, second row).

### v2

@@ Abstract
Introduction.– Simulating the non-Markovian dynamics of open quantum systems linearly coupled to Gaussian environments is a central challenge in quantum science [1, 2, 3, 4], particularly beyond the weak-coupling regime. When simulating the non-Markovian dynamics on a classical or quantum computer via an approximate representation, two key questions arise: (1) Efficiency: Can the simulation be performed using minimal computational resources? (2) Physicality: Does the approximate dynamics correspond to a physically realizable process?
We focus on approximation schemes that employ a finite number of auxiliary bath modes (often called the bath, environment, discrete modes, or pseudomodes in various settings). These auxiliary modes effectively capture the influence of the environment on the system dynamics. The efficiency of such schemes is characterized by the number of modes required to approximate the environment. The goal is to simulate all bounded system observables up to time $T$ within precision $\varepsilon$. This task is often reduced to reproducing the bath correlation functions (BCFs) with comparable precision [5, 6, 7, 8, 9]. The approximation scheme is considered efficient if the number of modes scales as $\text{polylog}(T/\varepsilon)$.
Physicality requires that the combined system-bath dynamics is formulated as a quantum channel, ensuring completely positive and trace-preserving (CPTP) dynamics. Lindbladian dynamics [10, 11] serves as a standard example of physical dynamics. The CPTP property, in turn, guarantees numerical stability when simulating the dynamics on a classical computer. Physicality is also essential for efficient implementation on a quantum computer [12, 13, 14, 15, 16, 17].
Table 1 summarizes and compares existing schemes. The most widely used approach is the unitary discrete mode representation, where both the system and auxiliary modes evolve under unitary dynamics [18, 19]. Despite its wide usage, the number of modes scales linearly in the simulation time $T$ [19, 20, 21]. Intuitively, this limitation arises because unitary dynamics of a finite system lack dissipation, making it incapable of accurately modeling BCFs that decay over time.
@@
Lorentzian pseudomode [22] $\mathrm{poly}(T/\varepsilon)$ ?
@@
The pseudomode theory [31, 32, 33, 5, 34, 27, 35, 36, 37, 38] aims at addressing this limitation by introducing dissipation in order to model energy relaxation correctly. In Refs. [31, 32, 5], the environment is represented by bath modes, each subject to Lindblad dissipation. In this scheme, the spectral density is expressed as a sum of Lorentzians, which we refer to as a Lorentzian pseudomode. However, a fundamental drawback is that the tail of a Lorentzian exhibits only inverse polynomial rather than exponential decay in the frequency domain. Theoretical analyses indicate that even for smooth spectral densities, the number of pseudomodes scales as $\text{poly}(T/\varepsilon)$ [22], though the preconstant can be smaller than that of unitary discrete modes.
@@
Recent advances in pseudomode methods, including non-Hermitian [23, 24, 39, 40, 41] and quasi-Lindblad pseudomodes [25, 26], have significantly improved the efficiency. Under certain analytic conditions on the spectral density, the required number of pseudomodes scales as $\mathrm{polylog}(T/\varepsilon)$ [7, 26]. However, these dynamics are not completely positive (CP) and cannot be realized as quantum channels.
@@
All the finite mode approximations discussed so far are decoupled, meaning that the modes interact with the system but not with one another. There is another class of approximation schemes that allows couplings between pseudomodes while preserving the Lindblad form [27, 28, 29, 30], and hence the resulting dynamics is physical. Although introduced in different contexts, we collectively refer to these as coupled Lindblad pseudomode theory. In particular, empirical studies in Refs. [27, 28] demonstrated that only a small number of such coupled pseudomodes are sufficient to approximate the BCF accurately. However, it remains unclear whether this provides a superior asymptotic scaling or simply a smaller prefactor. Moreover, the construction in [27, 28, 29, 30] relies on non-convex optimization, which can be challenging in practice.
@@
The contributions of this Letter are twofold: (1) We establish a direct connection between the coupled Lindblad and quasi-Lindblad pseudomode theory. Combined with the theoretical results of Ref. [26] on spectral density fitting, this work provides a theoretical justification for the efficiency of the coupled Lindblad pseudomodes with $\mathrm{polylog}(T/\varepsilon)$ scaling. (2) Inspired by the realization problem in control theory [42], we develop a robust numerical algorithm for constructing the coupled modes that avoids the non-convex optimization used in Refs. [27, 28, 29, 30]. This can significantly simplify the construction of coupled Lindblad pseudomodes, yielding higher accuracy while maintaining a minimal number of modes. Taken together, this provides a firm theoretical and computational foundation for the widespread application of pseudomode-based simulations.
@@
Model setup.– For concreteness, we consider a spin-boson Hamiltonian $\hat{H}=\hat{H}_{\text{S}}+\hat{H}_{\text{B}}+\hat{H}_{\text{SB}}$, where $\hat{H}_{\text{B}}=\int_{0}^{\infty}\omega\hat{b}_{\omega}^{\dagger}\hat{b}_{\omega}\mathrm{d}\omega$, $\hat{H}_{\text{SB}}=\hat{S}\hat{B}$, $\hat{S},\hat{B}$ are Hermitian, and $\hat{B}=\int_{0}^{\infty}\sqrt{J(\omega)}(\hat{b}_{\omega}+\hat{b}_{\omega}^{\dagger})\mathrm{d}\omega$. Here, $\hat{b}_{\omega}$ is a bosonic annihilation operator and $J(\omega)$ is the spectral density. Assuming an initially factorized state $\hat{\rho}(0)=\hat{\rho}_{\text{S}}(0)\otimes\hat{\rho}_{\text{B}}(0)$, where $\hat{\rho}_{\text{B}}(0)$ is the thermal equilibrium state of $\hat{H}_{\text{B}}$, our goal is to compute the dynamics of the system-reduced density operator $\hat{\rho}_{\text{S}}(t)=\operatorname{tr}_{\text{B}}\left({\mathrm{e}}^{-\mathrm{i}\hat{H}t}\hat{\rho}(0){\mathrm{e}}^{\mathrm{i}\hat{H}t}\right)$, where the influence of environment is fully captured by the BCF,
@@
Pseudomode theory.– Pseudomode theory introduces a finite number of auxiliary modes (denoted as A). If the pseudomode BCF matches the original BCF for all $t\in[0,T]$, then the corresponding pseudomode dynamics exactly reproduces the reduced system density operator $\hat{\rho}_{\text{S}}(t)$ up to time $T$ [5].
The dynamics of the density operator (denoted as $\hat{\rho}_{\text{SA}}^{\text{c}}(t)$) in the coupled Lindblad pseudomode is:
@@
with the bath initially in the vacuum state, i.e., $\hat{\rho}_{\text{A}}(0)=\mathinner{|{\boldsymbol{0}}\rangle}\!\!\mathinner{\langle{\boldsymbol{0}}|}$. Here, $H=H^{\dagger}$, $\Gamma\succeq 0$, and $\hat{A}=\sum_{k}g_{k}\hat{b}_{k}^{\dagger}+\overline{g_{k}}\hat{b}_{k}$, where we denote a positive semidefinite (definite) matrix $M$ as $M\succeq 0$ ($M\succ 0$). The conditions on $H$ and $\Gamma$ ensure that the dynamics of $\hat{\rho}_{\text{SA}}^{\text{c}}$ are CPTP, i.e. physical. The term coupled modes indicates that both $H$ and $\Gamma$ can be dense matrices; the off-diagonal elements mediate couplings between modes. In contrast, the Lorentzian pseudomode assumes $H$ and $\Gamma$ are diagonal, and the modes are decoupled. The BCF of the coupled Lindblad pseudomode, $C^{\text{c}}(t)=\operatorname{tr}\left(\hat{A}(t)\hat{A}(0)\hat{\rho}_{\text{A}}(0)\right)$, is given by [9, 28]:
@@
Another pseudomode theory in comparison is the quasi-Lindblad theory [25]. It includes additional system-bath dissipation to Eq. (2),
@@
The system-bath dissipation $\boldsymbol{D}_{\text{SA}}$ in general breaks the CP condition when there is no dissipation acting on the system. Violating the CP condition can induce instabilities in the quasi-Lindblad dynamics, posing challenges for classical simulation  [^1] There is a subtle concept called the Hamiltonian-induced stability [25], which may stabilize the quasi-Lindblad dynamics, even under the CP violation, but in the worst case, the quasi-Lindblad dynamics show the instability.. We note that the hierarchical equations of motion approach [47, 43] may also encounter similar stability challenges [48, 49, 50].
In the context of quantum simulation, violating the CP condition implies that the dynamics can no longer be efficiently  [^2] To encode the evolution using a quantum channel, an exponentially large subnormalization factor in $T$ must be introduced, rendering long-time simulation on a quantum computer prohibitively expensive. implemented as a quantum channel. In contrast, the coupled Lindblad dynamics are inherently CPTP and thus compatible with quantum hardware. The coupling between bosonic modes can also be realized in analog ion-trap-based quantum simulators [15, 52, 53, 16].
How do coupled modes achieve efficiency and physicality?– We first explain the physical intuition for the improved efficiency of coupled modes in Eq. 7. While the coupled modes offer greater expressive power via $\mathcal{O}(N^{2})$ parameters compared to the $\mathcal{O}(N)$ parameters of the decoupled pseudomode, it is not clear a priori whether it translates into a scaling advantage. Furthermore, the higher dimensionality of the parameter space makes the search for optimal parameters more computationally demanding.
We resolve such difficulties by establishing a connection between the coupled Lindblad and quasi-Lindblad pseudomode theory, the latter of which is known to achieve the $\mathrm{polylog}(T/\varepsilon)$ scaling. This connection is characterized by the feasibility condition in Eq. (7) of Eq. 7, obtained from the Hermiticity and complete positivity of the pseudomode dynamics. The feasibility condition also naturally leads to a robust numerical algorithm for constructing the pseudomodes using semidefinite programming (Eq. (8)).
Theoretical analysis for the $\mathrm{polylog}(T/\varepsilon)$ scaling and robust construction algorithms.–
We first discuss the connection between the coupled Lindblad BCF $C^{\text{c}}(t)=g^{\dagger}\mathrm{e}^{-\mathrm{i}Kt}g$ and the quasi-Lindblad BCF $C^{\text{q}}(t)=l^{\dagger}\mathrm{e}^{-\mathrm{i}\Lambda t}r$. This is achieved by introducing a gauge transformation that generates couplings between different modes, while enforcing Hermiticity and positivity. The result is summarized in the following theorem.
@@
$\exists\,Y\succ 0\ \text{s.t.}$ $Yr=l,\qquad$ (Hermiticity) (7)
$\mathrm{i}\!\left(Y\Lambda-\Lambda^{\dagger}Y\right)\succeq 0,\qquad$ (Positivity)
@@
The first part of the theorem parallels the result of Ref. [5], which establishes that, for fixed $\hat{H}_{\text{S}}$, $\hat{S}$, and $\hat{\rho}_{\text{S}}(0)$, the BCF uniquely determines the reduced system dynamics. Therefore, we focus on the second part. A gauge transformation, $\Lambda\rightarrow K=X\Lambda X^{-1}$, $l^{\dagger}\rightarrow l^{\dagger}X^{-1}$, $r\rightarrow Xr$, with an invertible matrix $X$, leaves $C^{\text{q}}(t)$ invariant. The gauge-transformed BCF takes the coupled Lindblad form if the following conditions hold: (a) $g=(l^{\dagger}X^{-1})^{\dagger}=Xr$, (b) $\Gamma=(K^{\dagger}-K)/2\mathrm{i}\succeq 0$. These two conditions correspond to the Hermiticity and the positivity of the dynamics, respectively. Introducing $Y=X^{\dagger}X$, we rewrite these as the equality and inequality constraints in Eq. 7 by multiplying $X$ and $X^{{\dagger}}$ to conditions (a) and (b). In addition, $X$ being invertible indicates that $Y\succ 0$, which makes conditions (a) and (b) equivalent to Eq. 7. ∎
Eq. 7 implies that when the feasibility condition is satisfied, the number of coupled Lindblad pseudomodes is no greater than that of quasi-Lindblad pseudomodes. Consequently, the $\mathrm{polylog}(T/\varepsilon)$ scaling of the latter extends directly to the coupled Lindblad pseudomode.
In practice, the feasibility condition may not hold exactly. Nonetheless, we construct a numerical procedure that minimally violates it by solving the following least-squares problem with semidefinite constraints:
@@
This problem could be solved efficiently via a semidefinite programming (SDP) solver. Setting $X=\sqrt{Y}$, we recover the parameters $g$, $H$, and $\Gamma$ in Eq. (3). This approach avoids the non-convex optimization used in prior works [28, 27, 29, 30]. We also propose a new procedure for obtaining coupled Lindblad parameters directly from the BCF in the frequency domain, motivated by the realization problem in control theory [54, 42], as explained in the Supplemental Material (SM) [55, Section S1].
General applicability.– The above discussion focuses on the single-site spin-boson model. Our coupled Lindblad pseudomode theory can be directly extended to multi-site systems as well as to fermionic environments. We refer to the SM [55], Section S2 and Section S3, for details.
@@
Figure 1: (a) For a fixed precision $\varepsilon=10^{-6}$ in fitting $C(t)$ for $t\in[0,T]$, we plot the number of modes required, NN, against the maximum simulation time $T$. The number of coupled Lindblad and quasi-Lindblad pseudomodes scales as $\mathcal{O}(\log T)$ in contrast to the $\mathcal{O}(T)$ scaling in the unitary and Lorentzian modes. (b) For a fixed $T=10$, we plot $\varepsilon$ versus NN, where the coupled Lindblad and quasi-Lindblad methods achieve a significantly faster convergence rate.
@@
Numerical results.– We begin by approximating the BCF $C(t)$ with coupled Lindblad modes to demonstrate their $\mathrm{polylog}(T/\varepsilon)$ scaling in comparison with other pseudomode approaches. Our target is to fit $C(t)$ for $t\in[0,T]$, derived from the Ohmic spectral density $J(\omega)=\omega\mathrm{e}^{-\omega/\omega_{c}}$ for $\omega\geq 0$ at zero temperature and $\omega_{c}=1$. For the coupled Lindblad fitting, we first fit $\widetilde{C}(\omega)$ in the frequency domain using the realization-based method, and further refine the result by using it as an initial guess for a gradient-based optimization of $C(t)$ in the time domain. We illustrate $T$-dependence (Fig. 1(a)) and $\varepsilon$-dependence (Fig. 1(b)) of the number of modes $N$, where we use the averaged $L^{2}$ error $\varepsilon$ [55, Eq. (S9)] as a measure of precision. See SM [55, Section S6] for more details on the numerical setup. The results of Fig. 1(a) with a fixed precision $\varepsilon=10^{-6}$ confirm that the coupled Lindblad pseudomode exhibits $N=\mathcal{O}(\log T)$ scaling, similar to the quasi-Lindblad pseudomode [26], in contrast to $N=\mathcal{O}(T)$ scaling in the unitary and Lorentzian modes. In Fig. 1(b), with a fixed $T=10$, all methods exhibit $N=\mathcal{O}(\log(1/\varepsilon))$ scaling, but the coupled Lindblad and quasi-Lindblad approaches converge significantly faster. The performance of the coupled Lindblad pseudomode closely matches that of the quasi-Lindblad pseudomode, indicating that the violation of the feasibility condition is small. Furthermore, we confirm that the $\mathrm{polylog}(T/\varepsilon)$ efficiency of the coupled pseudomodes remains valid across additional numerical examples, specifically considering the BCF from sub-Ohmic and semicircular spectral densities (see SM, [55, Fig. S1 and Fig. S3]).
@@
Next, we demonstrate that the coupled Lindblad pseudomode theory can accurately capture real-time system dynamics. We study the spin-boson model with $\hat{H}_{\text{S}}=\frac{\omega_{e}}{2}\hat{\sigma}_{z}$ and $\hat{S}=\hat{\sigma}_{x}$, where $\hat{\sigma}_{x/z}$ are Pauli operators. We follow the setup in Ref. [28] by choosing the Lorentzian-like spectral density, $J(\omega)=\frac{2g^{2}\kappa\omega_{c}\omega/\pi}{(\omega_{c}^{2}-\omega^{2})^{2}+\kappa^{2}\omega^{2}}$ for $\omega\geq 0$ and $J(\omega)=0$ for $\omega<0$. We focus on the ultra-strong coupling regime, with parameters $\omega_{c}=\omega_{e}=0.58,\ g=0.25,$ and $\kappa=0.1\text{meV}$ [28]. Fig. 2(a) describes population dynamics $n_{0}(t)=\langle 0|\hat{\rho}_{S}(t)|0\rangle$ evolved from the initial state $\hat{\rho}_{S}(0)=|0\rangle\!\langle 0|$. We use $N=2$ and $N=4$ coupled pseudomodes, compared to the reference data obtained from unitary dynamics with a large discretization $N=400$. We use a time-dependent density matrix renormalization group algorithm for numerical simulations (see SM [55, Section S4] for details). Our approach achieves accuracy comparable to the coupled pseudomode dynamics in Ref. [28] with only $N=4$ modes, versus their $N=10$. Notably, while Ref. [28] requires non-convex optimization and explicit penalties to suppress unphysical negative-frequency contributions, our construction inherently minimizes these contributions. In Fig. 2(b), the fitted BCF yields a minimal negative-frequency contribution of approximately $10^{-5}$ for $N=4$, which further vanishes to $10^{-9}$ for $N=10$. We also tested the method for the fermionic Anderson impurity model (see SM [55, Section S5]), demonstrating its applicability to fermionic environments.
@@
Figure 3: Normalized absorption spectrum $S(\omega)$ for the dimer model with two different environments $J_{0}(\omega)$ (a, c) and $J_{1}(\omega)$ (b, d), at zero (a, b) and finite temperature (77K, (c, d)).